We assume given a plane u passing by the tip of the vectors u=<-1,1,2, v=<2, - 1,0> and w=<1, 1, 3.

a.) Find the dot products u.v and w.v

b.) Determine whether or not there is a vector n that is perpendicular to u. If yes, then find the vector n. Otherwise explain why such a vector does not exist?


1
Expert's answer
2021-06-14T15:37:43-0400

a)


"\vec u\cdot \vec v=-1(2)+1(-1)+2(0)=-3"

"\vec w\cdot \vec v=1(2)+1(-1)+3(0)=1"

b)


"\vec u\times \vec v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 -1 & 1 & 2 \\
 2 & -1 & 0
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 1 & 2 \\
 -1 & 0
\end{vmatrix}-\vec j\begin{vmatrix}
 -1 & 2 \\
 2 & 0
\end{vmatrix}+\vec k\begin{vmatrix}
 -1 & 1 \\
 2 & -1
\end{vmatrix}"

"=-2\vec i+4\vec j-\vec k"

The vector "\vec n=-2\vec i+4\vec j-\vec k" is perpendicular to the vector "\vec u."


Find the equation of the plane "U"


"\begin{vmatrix}
 x-(-1) & y-1 & z-2 \\
 2-(-1) & -1-1 & 0-2 \\
 1-(-1) & 1-1 & 3-2 \\
\end{vmatrix}=0"

"(x+1)\begin{vmatrix}
 -2 & -2 \\
 0 & 1
\end{vmatrix}-(y-1)\begin{vmatrix}
 3 & -2 \\
 2 & 1
\end{vmatrix}+(z-2)\begin{vmatrix}
 3 & -2 \\
 2 & 0
\end{vmatrix}=0"

"-2x-2-7y+7+4z-8=0"

"2x+7y-4z+3=0"

The vector "\vec n=2\vec i+7\vec j-4\vec k" is perpendicular to the plane "U."